A father is 35 years old and his son is 2 years old. In how many years will the father be four times as old as his son??
Answer by stanbon(75887)

A father is 35 years old and his son is 2 years old. In how many years will the father be four times as old as his son??
-----
Let the number of years be "x":
Equation:
35+x = 2(2+x)
35+x = 4 + 2x
x = 31
------
Ans: 31 years
===========================
Cheers,
Stan H.
============